MARC (marc_d):

The expression @ganeshie8

MARC (marc_d):

\[ax^3-8x^2+bx+6\]is divisible by\[x^2-2x-3\]

MARC (marc_d):

Find the values of a and b.

ganeshie8 (ganeshie8):

Hint : factor \(x^2-2x-3\)

MARC (marc_d):

\[(x-1)^2\]

ganeshie8 (ganeshie8):

doesn't look right, try again

MARC (marc_d):

okay.

MARC (marc_d):

oh ya... xD

MARC (marc_d):

(X-3)(X+1)

ganeshie8 (ganeshie8):

That means the given polynomial is divisible by both the factors `x-3` and `x+1`

ganeshie8 (ganeshie8):

Remember the factor/remainder theorem ?

ganeshie8 (ganeshie8):

It looks something like this : If `x-k` is a factor of `f(x)`, then `f(k) = 0`

MARC (marc_d):

haven't learn yet. But i know how to use the long division method

ganeshie8 (ganeshie8):

Since `x-3` and `x+1` are the factors of given polynomial, if you plug x = 3 or x = -1 into the given polynomial, you should get 0.

ganeshie8 (ganeshie8):

\(ax^3-8x^2+bx+6\)

ganeshie8 (ganeshie8):

plug in x = 3 and set it equal to 0 plug in x = -1 and set it equal to 0

MARC (marc_d):

ookay,let me try.

MARC (marc_d):

9a-b=22 -a-b=2

MARC (marc_d):

is it correct?

ganeshie8 (ganeshie8):

I'm getting first equation as 9a + b = 22

ganeshie8 (ganeshie8):

second equation looks good

MARC (marc_d):

sorry,it ws typo error.

ganeshie8 (ganeshie8):

Ohk good :) two equations and two unknowns, you can solve them

MARC (marc_d):

i got a=3 and b=-5

MARC (marc_d):

Thank you! @ganeshie8 ^_^